# supplementary

### v1

@@ Abstract
Recent advancements in digital pathology have enabled comprehensive analysis of Whole-Slide Images (WSI) from tissue samples, leveraging high-resolution microscopy and computational capabilities. Despite this progress, there is a lack of labeled datasets and open source pipelines specifically tailored for analysis of skin tissue. Here we propose Histo-Miner, a deep learning-based pipeline for analysis of skin WSIs and generate two datasets with labeled nuclei and tumor regions. We develop our pipeline for the analysis of patient samples of cutaneous squamous cell carcinoma (cSCC), a frequent non-melanoma skin cancer. Utilizing the two datasets, comprising 47,392 annotated cell nuclei and 144 tumor-segmented WSIs respectively, both from cSCC patients, Histo-Miner employs convolutional neural networks and vision transformers for nucleus segmentation and classification as well as tumor region segmentation. Performance of trained models positively compares to state of the art with multi-class Panoptic Quality (mPQ) of 0.569 for nucleus segmentation, macro-averaged F1 of 0.832 for nucleus classification and mean Intersection over Union (mIoU) of 0.884 for tumor region segmentation. From these predictions we generate a compact feature vector summarizing tissue morphology and cellular interactions, which can be used for various downstream tasks. Here, we use Histo-Miner to predict cSCC patient response to immunotherapy based on pre-treatment WSIs from 45 patients. Histo-Miner predicts patient response with mean area under ROC curve of 0.755 $\pm$ 0.091 over cross-validation, and identifies percentages of lymphocytes, the granulocyte to lymphocyte ratio in tumor vicinity and the distances between granulocytes and plasma cells in tumors as predictive features for therapy response. This highlights the applicability of Histo-Miner to clinically relevant scenarios, providing direct interpretation of the classification and insights into the underlying biology. Importantly, Histo-Miner is designed to allow for its use on other cancer types and on other training datasets. Our tool and datasets are available through our github repository: https://github.com/bozeklab/Histo-Miner.
@@ 2.1 Histo-Miner pipeline description
After determining tumor areas using SCC Segmenter, the results of the SCC Hovernet cell nuclei classification are updated to add a new cell class as follows: all the nuclei predicted as tumor cells outside of the predicted tumor regions are reclassified as epithelial. The reason for this update is that healthy epithelial cells and tumor cells have similar morphologies and are impossible to discriminate without a broader context and information about the tissue structure. Example visualization of the inference of the two methods is shown in Fig. 2.
@@ 2.2 Nuclei segmentation and classification evaluation
To additionally validate Histo-Miner, we performed immunohistochemistry (IHC) for Mypo (Granulocytes), CD3 (lymphocytes), CD79a (plasma cells), CD10 (stromal cells) and p40 (tumor cells), as well as H&E staining on adjacent slides from the same tissue blocks. In 6 different cSCC tumors we selected 7 to 11 representative ROIs (750µm x 750µm) on H&E-stained image and predicted the number of cells of each type using Histo-Miner (Fig. 3c). The same regions in IHC images were classified by a board certified dermatopathologist into 4 groups (- , +, ++, $+$) based on the level of IHC positivity. Comparing the number of predicted cells across IHC positivity groups, we observed a significant association of the number of predicted cells per cell type and IHC positivity of the appropriate marker (Fig. 3d).
@@ 2.4 Application of Histo-Miner to predict response of cSCC patients to immunotherapy
We collected WSIs from 45 patients (one per patient) with cSCC skin cancer, from 6 medical centers before they received immune checkpoint inhibitors treatment. 28 of them were classified as responders, showing partial response (PR) or complete response (CR), i.e. tumor shrinkage. 17 patients were classified as non-responders, showing stable disease (SD) or progressive disease (PD) states (Fig. 4a). This collection of classified WSIs represents our fourth dataset, called CPI, (openly available, see Data availability section).
We used Histo-Miner to extract tissue representative features from the WSIs of CPI dataset. Then, we trained and evaluated an XGBoost classifier [24] for the task of classifying patients in their two categories based on the feature vectors. We evaluated XGBoost classifier through 3 fold cross-validation and performed feature selection within the cross-validation runs. We used minimum Redundancy - Maximum Relevance (mRMR) feature selection method [25] which is designed to find the smallest relevant subset of features (maximum relevance) while preventing highly correlated features to be part of this subset (minimum redundancy). All the 107 features kept for analysis are describing tissue structure, no nucleus morphology features (area, circularity) were included (See Tissue Analyser section in Methods for an in-depth description of the features).
Due to the limited number of samples in the dataset (45) we could not perform hyperparameter search within nested cross-validation so we kept the default set of hyperparameters for all the cross-validation runs. In each of the runs, we performed feature selection with mRMR and ranked all features from most to least predictive. We then trained 107 XGBoost classifiers on the same training split, iteratively removing the least predictive feature each time, thus going from all 107 features down to the single most predictive feature. For each of these trainings we calculated balanced accuracy on the validation split. To know how many features to keep we calculated the average balanced accuracy for each training keeping $N\in[1,107]$ features. We kept $N_{\text{best}}=19$ features which corresponded to the best mean balanced accuracy across all runs. We obtained mean balanced accuracy $\bar{ba}_{N_{\text{best}}}=0.767\pm 0.057$ across 3 fold cross-validation runs with the mean area under ROC curve $\bar{AUC}_{N_{\text{best}}}=0.755\pm 0.091$ (Fig. 4b ). Interestingly, adding the morphology features to the set of features did not improve the classification - with $N_{\text{best}_{\text{w/morph}}}=305$ features the balanced accuracy reaches $\bar{ba}_{N_{\text{best}_{\text{w/morph}}}}=0.754\pm 0.127$.
The 4 most representative features in the decreasing ranking order are: the percentage of lymphocytes among all cells in tumor vicinity, the ratio between the number of granulocytes and lymphocytes in tumor vicinity, the percentage of lymphocytes among all cells in tumor region, and the mean distances between granulocytes and plasma cells in tumor regions (Fig. 4). We discuss the interpretation of these features in the next section. The process behind feature selection is described in the Method section.
@@ 3 Discussion
4 Methods
4.1 NucSeg and TumSeg datasets descriptions
To build a tumor segmenter algorithm, we additionally assembled 144 WSIs of 125 cSCC patients from 3 medical centers. These WSIs are a subset of the dataset described in [39]. All tumor regions were labeled by two pathology experts. The WSIs were originally at resolution 40x and downsampled 32 times (reaching a final resolution of 1.25x) to enable to fit them into memory during model training (Fig. 5a). Examples of WSI tumor segmentation annotations are displayed in Fig. 5c. WSIs and the binary segmentation masks of their tumor regions are part of TumSeg dataset.
4.2 Histo-Miner deep learning models
Histo-Miner implementation includes deep learning models trained with our custom datasets and hyperparameter values determined using a grid-search strategy. Histo-Miner users can utilize the weights of these models to perform similar inferences on their own datasets, re-train these models through Histo-Miner implementation directly and edit model architectures and hyperparameters for further development. To fit user needs, it is possible to use only specific blocks of the pipeline - such as inference of the deep learning models - instead of using the whole process until calculation of tissue features.
4.3 Training SCC Hovernet
We trained NucSeg dataset including 5,968 patches of size 540x540 pixels with 70% overlap. Our validation set contained 848 patches of size 540x540 pixels with 70% overlap. We tested 3 different training and validation batch sizes - 4, 8, 16 - and 3 different Adam decay rates for momentum estimation (Adam betas) - 0.8, 0.9, 0.999. We tested different training strategies combining network pre-trained on ImageNet 21k, pre-trained on the not-curated H&E nucleus dataset, and trained from scratch. Performance of resulting models was highest when the model was first pre-trained on ImageNet 21k, then pre-trained on the not-curated H&E nucleus dataset before fine-tuning. The choice of final model was based on maximizing the panoptic segmentation task performance.
4.4 Training SCC Segmenter
We performed hyperparameter grid-search to find the best set of hyperparameters as described in Supplementary Table. 1.
4.5 Tissue Analyser
Within Histo-Miner, SCC Segmenter model segments tumor regions. SCC Hovernet model performs instance segmentation of cell nuclei and classifies them into different cell classes (granulocytes, lymphocytes, plasma cells, stromal cells, and tumor cells). Using both models’ predictions, we refine the cell classification and add one more class among the possible predictions, the epithelial class. In fact, healthy epithelial cells and tumor cells are hard to discriminate without a broader context and information about the tissue structure. Healthy epithelial cells and tumor cells have similar morphologies. To distinguish these two cell types, we added one refinement step: all the nuclei predicted as belonging to tumor cells by SCC Hovernet, located outside the tumor regions predicted by Segmenter, are reclassified as epithelial. The result of the classification update is visible in Fig. 2.
One of our features is the average distance of the closest cell of a given class X - source class - to the closest cell of a given class Y - target class - inside the tumor regions. The average distance is calculated as shown in Eq. 1 and through the following steps: 1- We generate a rectangle that defines the initial search area. The rectangle height is 5% of the tumor bounding box height and width 5% of the tumor bounding box width. It is centered around the nucleus of the source cell class. 2- We verify if there is at least one nucleus of the target cell class inside this search area. 3- If there is at least one, we calculate all distances between source and target cells and keep the smallest one. If there is no nucleus of target cell class we increase the search area until we find at least one nucleus of target cell class in the search area. The search area cannot extend to other tumor areas. 4- We perform steps 1-3 for all nuclei of the source class. A quantitative explanation is available in Algorithm. 1 (all the memory optimization steps are skipped for readability). The increase of search area for each iteration was optimized to reduce calculation time. For computation optimization reasons, the search area is a rectangle and not a circle. Indeed, one of the main reason is that searching for coordinates in a rectangle is faster than searching for coordinates in a circle (only comparisons instead of subtractions and multiplications). In some specific cases, using a rectangle search area can lead to overestimation of the distance. Description of these cases, probability of overestimation, and bounding of overestimation are described in Supplementary Fig. 2. This probability of overestimation decreases drastically with the number of cells in the tumor. For instance, we can calculate that for a squared search area of side length $2r_{1}$ and origin 00, if $N$ cells are in the circle of radius $\sqrt{2}r_{1}$ and origin 00 (the square is inscribed in this circle), the probability of overestimation is $P(error_{N})_{2}=0.022$ for $N=2$ and $P(error_{N})_{10}=2.8\mathrm{x}10^{-4}$ for $N=10$. These distances describe the interactions between different cell types inside the tissue. They are calculated for granulocytes, lymphocytes, plasma cells, and tumor cells to assess the organization of the tissue regarding the intensity of the immune response.
[…]
20
[…]
$\text{with}\ \lambda_{0}\ \textrm{verifying:}$:
[…]
{adjustwidth}
30pt30pt where $\lambda\in\mathcal{N}^{*}$ coefficient of increase of sides length of search rectangle, $l_{t}$ and $w_{t}$ length and width of the bounding box of the tumor region, $E_{c_{a}}$ set of coordinates in $\mathcal{N}^{2}$ of all cells centroid of class A (source class), $E_{c_{b}}$ set of coordinates in $\mathcal{N}^{2}$ of all cells centroid of class B (target class), $n_{c_{A}}$ number of cells of class A.
Algorithm 1 Minimum Distance Calculation Pseudo-Code
Require: $classjson$: file with centroid and type of all WSI’s cells
Require: $maskmap$: segmentation of WSI’s tumor regions
Require: $selected\_classes$: classes for which we want to calculate distances, $f$: acceleration of increase parameter - here $f=0.05$
In each process a different pair source class / target class is defined
For each process:
tab $allmindist$ = list()
tab For all cells of source class:
tabtab $dist\_list$ = list()
tabtab Check tumor ID of source cell
tabtab List all target class cells $centroid$ $coordinates$ in the same Tumor
tabtab $\lambda=1$
tabtab $kept\_targets$ = list()
tabtab While $kept\_targets$ cells list is empty:
tabtabtab Append to $kept\_targets$ all target cells $centroid$ $coordinates$ inside the search area
tabtabtab $\lambda=\lambda+1$
tabtab For all cell in $kept\_targets$:
tabtabtab $(x_{k},y_{k})$ coordinates of cells of target class
tabtabtab $dist\_list\mathbf{.append}(dist)$
tabtab $min\_dist$ = min($dist\_list$)
tabtzb $allmindist$.append($min\_dist$)
Ouput a list of all $avgdist$ gathered items
[…]
We performed 3 fold cross-validation with XGBoost classifier [24] on the CPI dataset to predict the class of the WSIs.
The CPI dataset consists of 45 WSIs from cSCC skin cancer of 6 medical centers before they received immune checkpoint inhibitors treatment. Each WSI is taken from a different patient. 28 of them were classified as responders (CR and PR states) and 17 patients were classified as non-responders (SD and PD states) as displayed in Fig. 4a.
The low number of folds is constrained by the low number of samples. In fact, having too small validation folds would increase variance in evaluation. We kept $N_{best}=19$ features, selected using mRMR method [25], which corresponded to the best mean balanced accuracy across all runs as described in Results section.
Even if the number of selected features $N_{best}$ is the same for each run in the case of mRMR, the selection of features can vary. Following Eq. 2, we identify the most representative features by selecting those with the highest occurrence counts ${c_{f_{k}}}$ across the selected feature sets in each cross-validation run. In the event of ties, we favor features with higher rankings in their respective sets. To do so we first record the position of the feature in its set, its rank, and calculate its pre-score as $10^{N_{best}-rank}$, so a feature ranked first would have the highest pre-score. Then we add all the pre-scores for each cross-validation groups for a given feature and take the log of this sum to obtain the final score $s_{f_{k}}$.
[…]
10
$c_{f_{k}}=\sum_{l=1}^{L}\sum_{n=1}^{N_{b}}\delta(A_{l,n},f_{k})$
$s_{{f_{k}}}=\log(\sum_{l=1}^{L}10^{{N_{best}-rank(f_{k})_{split_{l}}}})$
{adjustwidth}
30pt30pt where $fv_{split_{l}}$ = $[f_{1},f_{2},...f_{N_{b}}]_{split_{l}}$ the vector of $N_{best}$ selected features from split $l\in[1,L]$ of the cross-validation, $rank(f_{k})_{split_{l}}$ the rank of feature $f_{k}$ selected from split $l\in[1,L]$ and $\delta\text{ the Kronecker delta}$.
Other feature selection methods were tested but lead to a poorer selection of features. We tried Boruta [44] method. Unlike mRMR that is designed to find the smallest relevant subset of features, Boruta tries to find all features carrying information usable for prediction. With Boruta the number of selected features varied for each fold. Classification of CPI WSIs using Boruta selected feature had a balanced accuracy of $\bar{ba}_{N_{{BORUTA}}}=\bar{ba}_{[5,0,1]}=0.711\pm 0.011$ across the 3 folds. For one of the fold Boruta was enable to select features, and this for any setting of maximum depth of the random tree. In the best scenario presented here the maximum depth was set to 1, the random state was set to 0 and all other hyperparameters were kept as default values. We also ranked features following Mann-Whitney U test scores [45]. We then trained 107 XGBoost classifiers on the same training split, iteratively removing the least predictive feature each time, thus going from all 107 features down to the single most predictive feature (as we did for mRMR). The best balanced accuracy was in this case $\bar{ba}_{N_{MannWhitney}}=0.715\pm 0.069$ with $N_{MannWhitney}=6$.
@@ 4.6 Feature selection to predict response of cSCC patients to immunotherapy
Finally the CPI image classification dataset utilized as a use-case for Histo-Miner is also available on Zenodo repository link. All the WSIs of these datasets were annonymized and cannot be used for commercial use.

### v2

@@ Abstract
Recent advances in digital pathology have enabled comprehensive analyses of Whole-Slide Images (WSIs) from tissue samples, leveraging high-resolution microscopy and computational capabilities. Despite this progress, available tools for automatic cell type identification perform poorly on skin tissue, e.g. in the classification of non-melanoma tumor cells. This is due to a paucity of labeled training data sets and high morphological similarities between tumor and non-tumor epithelial cells in the skin. Here, we propose Histo-Miner, a deep learning-based pipeline designed for the analysis of skin WSIs. To this end we generated two new datasets using WSIs of cutaneous Squamous Cell Carcinoma (cSCC) samples, a frequent non-melanoma skin cancer, by annotating 47,392 cell nuclei across 5 cell types in 21 WSIs and segmenting tumor regions in 144 WSIs. Histo-Miner employs convolutional neural networks and vision transformers for nucleus segmentation and classification, as well as tumor region segmentation. Performance of trained models positively compares to state of the art with multi-class Panoptic Quality (mPQ) of 0.569 for nucleus segmentation, macro-averaged F1 of 0.832 for nucleus classification and mean Intersection over Union (mIoU) of 0.884 for tumor region segmentation. From these output, the pipeline can generate a compact feature vector summarizing tissue morphology and cellular interactions, which can be used for various downstream tasks. As an exemplary use-case, we deploy Histo-Miner to predict cSCC patient response to immunotherapy based on pre-treatment WSIs from 45 patients. Histo-Miner predicts patient response with mean area under ROC curve of 0.755 $\pm$ 0.091 over cross-validation, and identifies percentages of lymphocytes, the granulocyte to lymphocyte ratio in tumor vicinity and the distances between granulocytes and plasma cells in tumors as predictive features for therapy response. This highlights the applicability of Histo-Miner to clinically relevant scenarios, providing direct interpretation of the classification and insights into the underlying biology. Importantly, Histo-Miner is designed to allow for its use on other cancer types and on other training datasets. Our tool and datasets are available through our github repository: https://github.com/bozeklab/histo-miner.
@@ 2.1 Histo-Miner pipeline description
After determining tumor areas using SCC Segmenter, the results of the SCC Hovernet cell nuclei classification are updated to add a new cell class as follows: all the nuclei predicted as tumor cells outside of the predicted tumor regions are reclassified as healthy epithelial. The reason for this update is that healthy epithelial cells and tumor cells have similar morphologies and are impossible to discriminate without a broader context and information about the tissue structure (see Supplementary Fig.1). Example visualization of the inference of the two methods is shown in Fig. 2.
@@
2.2 NucSeg and TumSeg datasets descriptions
To build a tumor segmenter algorithm, we additionally assembled 144 WSIs of 125 cSCC patients from 3 medical centers - Bonn, Cologne and Munich. These WSIs are a subset of the dataset described in [19]. All tumor regions were labeled by two pathology experts. The WSIs were originally at resolution 40x and downsampled 32 times (reaching a final resolution of 1.25x) to enable to fit them into memory during model training (Fig. 3a). Examples of WSI tumor segmentation annotations are displayed in Fig. 3c. WSIs and the binary segmentation masks of their tumor regions are part of TumSeg dataset. In addition, anonymized patient IDs and medical centers IDs are available as metadata.
2.3 Histo-Miner deep learning models
Histo-Miner implementation includes deep learning models trained with our custom datasets. Histo-Miner users can utilize the weights of these models to perform similar inferences on their own datasets, re-train these models through Histo-Miner implementation directly and edit model architectures and hyperparameters for further development. To fit user needs, it is possible to use only specific blocks of the pipeline - such as inference of the deep learning models - instead of using the whole process until calculation of tissue features.
2.4 Training SCC Hovernet
We trained NucSeg dataset including 5,968 patches of size 540x540 pixels with 70% overlap. Our validation set contained 848 patches of size 540x540 pixels with 70% overlap. We kept the same set of hyperparameters as the original implementation [14], except that we doubled the training batch size for our last training step of 200 epochs (see above). We tested different training strategies combining network pre-trained on ImageNet 21k, pre-trained on the not-curated H&E nucleus dataset, and trained from scratch. Performance of resulting models was highest when the model was first pre-trained on ImageNet 21k, then pre-trained on the not-curated H&E nucleus dataset before fine-tuning. The choice of final model was based on maximizing the panoptic segmentation task performance.
2.5 Training SCC Segmenter
The model was trained on randomly chosen 115 slides of TumSeg dataset and the validation set contained the remaining 29 slides of the dataset. We performed hyperparameter grid-search to find the best set of hyperparameters as described in Supplementary Table. 1. The accuracy estimation was performed on the validation set due to the limited number of slides and lack of an independent test set.
2.6 Tissue Analyser
Within Histo-Miner, SCC Segmenter model segments tumor regions. SCC Hovernet model performs instance segmentation of cell nuclei and classifies them into different cell classes (granulocytes, lymphocytes, plasma cells, stromal cells, and tumor cells). Using both models’ predictions, we refine the cell classification and add one more class among the possible predictions, the healthy epithelial class. In fact, healthy epithelial cells and tumor cells are hard to discriminate without a broader context and information about the tissue structure. Healthy epithelial cells and tumor cells have similar morphologies. To distinguish these two cell types, we added one refinement step: all the nuclei predicted as belonging to tumor cells by SCC Hovernet, located outside the tumor regions predicted by Segmenter, are reclassified as epithelial. The result of the classification update is visible in Fig. 2.
One of our features is the average distance of the closest cell of a given class X - source class - to the closest cell of a given class Y - target class - inside the tumor regions. The average distance is calculated as shown in Eq. 1 and through the following steps: 1- We generate a rectangle that defines the initial search area. The rectangle height is 5% of the tumor bounding box height and width 5% of the tumor bounding box width. It is centered around the nucleus of the source cell class. 2- We verify if there is at least one nucleus of the target cell class inside this search area. 3- If there is at least one, we calculate all distances between source and target cells and keep the smallest one. If there is no nucleus of target cell class we increase the search area until we find at least one nucleus of target cell class in the search area. The search area cannot extend to other tumor areas. 4- We perform steps 1-3 for all nuclei of the source class. A quantitative explanation is available in Algorithm. 1 (all the memory optimization steps are skipped for readability). The increase of search area for each iteration was optimized to reduce calculation time. For computation optimization reasons, the search area is a rectangle and not a circle. Indeed, one of the main reason is that searching for coordinates in a rectangle is faster than searching for coordinates in a circle (only comparisons instead of subtractions and multiplications). In some specific cases, using a rectangle search area can lead to overestimation of the distance. Description of these cases, probability of overestimation, and bounding of overestimation are described in Supplementary Fig. 3. This probability of overestimation decreases drastically with the number of cells in the tumor. For instance, we can calculate that for a squared search area of side length $2r_{1}$ and origin 0, if $N$ cells are in the circle of radius $\sqrt{2}r_{1}$ and origin 0 (the square is inscribed in this circle), the probability of overestimation is $P(error_{N})_{2}=0.022$ for $N=2$ and $P(error_{N})_{10}=2.8\mathrm{x}10^{-4}$ for $N=10$. These distances describe the interactions between different cell types inside the tissue. They are calculated for granulocytes, lymphocytes, plasma cells, and tumor cells to assess the organization of the tissue regarding the intensity of the immune response.
[…]
$\text{with}\ \lambda_{0}\ \textrm{verifying:}$
[…]
where $\lambda\in\mathcal{N}^{*}$ coefficient of increase of sides length of search rectangle, $l_{t}$ and $w_{t}$ length and width of the bounding box of the tumor region, $E_{c_{a}}$ set of coordinates in $\mathcal{N}^{2}$ of all cells centroid of class A (source class), $E_{c_{b}}$ set of coordinates in $\mathcal{N}^{2}$ of all cells centroid of class B (target class), $n_{c_{A}}$ number of cells of class A.
Algorithm 1 Minimum Distance Calculation Pseudo-Code
Require: classjsonclassjson: file with centroid and type of all WSI’s cells
Require: maskmapmaskmap: segmentation of WSI’s tumor regions
Require: selected_classesselected\_classes: classes for which we want to calculate distances, $f$: acceleration of increase parameter - here $f=0.05$
In each process a different pair source class / target class is defined
For each process:
tab allmindistallmindist = list()
tab For all cells of source class:
tabtab dist_listdist\_list = list()
tabtab Check tumor ID of source cell
tabtab List all target class cells centroidcentroid coordinatescoordinates in the same Tumor
tabtab $\lambda=1$
tabtab kept_targetskept\_targets = list()
tabtab While kept_targetskept\_targets cells list is empty:
tabtabtab Append to kept_targetskept\_targets all target cells centroidcentroid coordinatescoordinates inside the search area
tabtabtab $\lambda=\lambda+1$
tabtab For all cell in kept_targetskept\_targets:
tabtabtab $(x_{k},y_{k})$ coordinates of cells of target class
tabtabtab $dist\_list\mathbf{.append}(dist)$
tabtab min_distmin\_dist = min(dist_listdist\_list)
tabtzb allmindistallmindist.append(min_distmin\_dist)
Ouput a list of all avgdistavgdist gathered items
Other features include ratios of cell types and percentages of cells of a given cell type in the vicinity of the tumor, in the tumor regions or in the whole WSI. The vicinity of tumor is defined as a $1\textrm{mm}$-wide area around the tumor [23]. In the case of ratios, to limit outlayers we define ratio $\eta=\frac{\log{(n_{c_{A}})}+\epsilon}{\log{(n_{c_{B}})}+\epsilon}$. with $n_{c_{A}}$ number of cells of class A, $n_{c_{B}}$ number of cells of class B and $\epsilon=10^{-3}$, smoothness parameter. The full list of features is in the Supplementary Data.
We collected WSIs from 45 patients (one per patient) with cSCC skin cancer, from 6 medical centers - Bonn, Cologne, Dortmund, Munich, Oberhausen and Salzburg - before they received immune checkpoint inhibitors treatment. 28 of them were classified as responders, showing partial response (PR) or complete response (CR), i.e. tumor shrinkage. 17 patients were classified as non-responders, showing stable disease (SD) or progressive disease (PD) states. More specifically, CR means disappearance of all lesions; PR: 50% or more in decrease of total tumor size; SD: <<50% decrease and/or <<25% increase of one/several tumor lesions; PD: >>25% increase of one/several tumor lesions or new lesions. The classification was determined by a dedicated review of the clinical and radiological imaging by at least two observations not less than 4 weeks apart (following World Health Organization handbook for reporting results of cancer treatment). This collection of classified WSIs represents our fourth dataset, called CPI (see Data availability section), and is succinctly described in Fig. 5a.
We used Histo-Miner to extract tissue representative features from the WSIs of CPI dataset. Then, we trained and evaluated an XGBoost classifier [24] for the task of classifying patients in their two categories based on the feature vectors. We evaluated XGBoost classifier through 3 fold cross-validation and performed feature selection within the cross-validation runs. Due to the limited number of samples in the dataset (45) we could not perform hyperparameter search within nested cross-validation so we kept the default set of hyperparameters for all the cross-validation runs. Similarly, the low number of folds is constrained by the low number of samples. In fact, having too small validation folds would increase variance in evaluation. We used minimum Redundancy - Maximum Relevance (mRMR) feature selection method [25] which is designed to find the smallest relevant subset of features (maximum relevance) while preventing highly correlated features to be part of this subset (minimum redundancy). All the 107 features kept for analysis are describing tissue structure, no nucleus morphology features (area, circularity) were included (See Tissue Analyser section in Methods for an in-depth description of the features).
To know how many features to keep we calculated the average balanced accuracy for each training keeping $N\in[1,107]$ features. We kept $N_{\text{best}}=19$ features which corresponded to the best mean balanced accuracy across all runs. Even if the number of selected features $N_{best}$ is the same for each run in the case of mRMR, the selection of features can vary. Following Eq. 2, we identify the most representative features by selecting those with the highest occurrence counts $c_{f_{k}}$} across the selected feature sets in each cross-validation run. In the event of ties, we favor features with higher rankings in their respective sets. To do so we first record the position of the feature in its set, its rank, and calculate its pre-score as $10^{N_{best}-rank}$, so a feature ranked first would have the highest pre-score. Then we add all the pre-scores for each cross-validation groups for a given feature and take the log of this sum to obtain the final score $s_{f_{k}}$.
$c_{f_{k}}=\sum_{l=1}^{L}\sum_{n=1}^{N_{b}}\delta(A_{l,n},f_{k})$
$s_{{f_{k}}}=\log(\sum_{l=1}^{L}10^{{N_{best}-rank(f_{k})_{split_{l}}}})$
where $fv_{split_{l}}$ = $f_{1},f_{2},...f_{N_{b}}]_{split_{l}}$ the vector of $N_{best}$ selected features from split $l\in[1,L]$ of the cross-validation, $rank(f_{k})_{split_{l}}$ the rank of feature $f_{k}$ selected from split $l\in[1,L]$ and $\delta\text{ the Kronecker delta}$.
Other feature selection methods were tested but led to a worse performing feature set. We tested Boruta [26] method which, unlike mRMR that finds the smallest relevant subset of features, aims at finding all features carrying information usable for prediction. With Boruta the number of selected features varied for each fold. Classification of CPI WSIs using Boruta selected feature had a balanced accuracy of $\bar{ba}_{N_{{BORUTA}}}=\bar{ba}_{[5,0,1]}=0.711\pm 0.011$ across the 3 folds. For one of the fold Boruta was enable to select features, and this for any setting of maximum depth of the random tree. In the best scenario presented here the maximum depth was set to 1, the random state was set to 0 and all other hyperparameters were kept as default values. We also ranked features following Mann-Whitney U test scores [27]. We then trained 107 XGBoost classifiers on the same training split, iteratively removing the least predictive feature each time, thus going from all 107 features down to the single most predictive feature (as we did for mRMR). The best balanced accuracy was in this case $\bar{ba}_{N_{MannWhitney}}=0.715\pm 0.069$ with $N_{MannWhitney}=6$.
3 Results
@@ 3.1 Nuclei segmentation and classification evaluation
To additionally validate Histo-Miner, we performed immunohistochemistry (IHC) for Mypo (Granulocytes), CD3 (lymphocytes), CD79a (plasma cells), CD10 (stromal cells) and p40 (tumor cells), as well as H&E staining on adjacent slides from the same tissue blocks. In 6 different cSCC tumors we selected 7 to 11 representative ROIs (750µm x 750µm) on H&E-stained image and predicted the number of cells of each type using Histo-Miner (Fig. 4c). The same regions in IHC images were classified by a board certified dermatopathologist into 4 groups (- , +, ++, +++) based on the level of IHC positivity. Comparing the number of predicted cells across IHC positivity groups, we observed a significant association of the number of predicted cells per cell type and IHC positivity of the appropriate marker (Fig. 4d).
@@ 3.3 Application of Histo-Miner to predict response of cSCC patients to immunotherapy
The CPI dataset consists of 45 WSIs from cSCC skin cancer of 6 medical centers before they received immune checkpoint inhibitors treatment. Each WSI is taken from a different patient. 28 of them were classified as responders (CR and PR states) and 17 patients were classified as non-responders (SD and PD states) as displayed in Fig. 5a. We performed 3 fold cross-validation with XGBoost classifier [24] on the CPI dataset to predict the class of the WSIs.
In each of the runs, we performed feature selection with mRMR, as described in Methods section, and ranked all features from most to least predictive. We then trained 107 XGBoost classifiers on the same training split, iteratively removing the least predictive feature each time, thus going from all 107 features down to the single most predictive feature. For each of these trainings we calculated balanced accuracy on the validation split. To know how many features to keep we calculated the average balanced accuracy for each training keeping $N\in[1,107]$ features. We kept $N_{\text{best}}=19$ features which corresponded to the best mean balanced accuracy across all runs. We obtained mean balanced accuracy $\bar{ba}_{N_{\text{best}}}=0.767\pm 0.057$ across 3 fold cross-validation runs with the mean area under ROC curve $\bar{AUC}_{N_{\text{best}}}=0.755\pm 0.091$ (Fig. 5b ). Interestingly, adding the morphology features to the set of features did not improve the classification - with $N_{\text{best}_{\text{w/morph}}}=305$ features the balanced accuracy reaches $\bar{ba}_{N_{\text{best}_{\text{w/morph}}}}=0.754\pm 0.127$.
The 4 most representative features in the decreasing ranking order are: the percentage of lymphocytes among all cells in tumor vicinity, the ratio between the number of granulocytes and lymphocytes in tumor vicinity, the percentage of lymphocytes among all cells in tumor region, and the mean distances between granulocytes and plasma cells in tumor regions (Fig. 5c). We discuss the interpretation of these features in the next section. The process behind feature selection is described in the Method section.
@@ 4 Discussion
Finally the CPI image classification dataset utilized as a use-case for Histo-Miner is also available on Zenodo repository link (following paper publication).
All the WSIs of these datasets were anonymized and cannot be used for commercial use.